OpenStudy (anonymous):

What is the equation of a line passing through (-6, 5) and having a slope of ?

OpenStudy (anonymous):

@Smallville133760

OpenStudy (anonymous):

Hey

OpenStudy (anonymous):

Give me your options

OpenStudy (johnweldon1993):

also make sure to include the missing slope from the question :)

OpenStudy (anonymous):

And that!

OpenStudy (anonymous):

y = 3x + 7

OpenStudy (anonymous):

y=1/3x-23/3 y=1/3x-3

OpenStudy (anonymous):

thoes are all the choices

OpenStudy (anonymous):

y = 3x + 7

OpenStudy (johnweldon1993):

What is the slope from the question? (it was omitted when you posted it)

OpenStudy (anonymous):

of 1/3

OpenStudy (anonymous):

hmm

OpenStudy (johnweldon1993):

Okay...so if the slope = 1/3 and the point is (-6,5) We can use a point slope form \[\large y - y_0 = m(x - x_0)\] and plug in all our numbers... \[\large y - 5 = \frac{1}{3}(x + 6)\] Can you distribute and simplify this?

OpenStudy (anonymous):

uhh i can try ;_:

OpenStudy (johnweldon1993):

go ahead I'll be here if you get lost :)

OpenStudy (anonymous):

ok so you multiplt 1/3 by x so you get 1/3x then 1/3 by 6 and get 6/3 wich simplifys to 2? then 1/3 X 2 and you 2/3 ...then what?

OpenStudy (anonymous):

nm the las t sentence thats wrong

OpenStudy (johnweldon1993):

Right the last sentence was off...but otherwise correct :) so now we have \[\large y - 5= \frac{1}{3}x + 2\] Now we want to solve for 'y' ..so we need to add 5 to both sides right?

OpenStudy (anonymous):

y=1/3x+7

OpenStudy (johnweldon1993):

And that's your answer

OpenStudy (anonymous):

thank could you help mee some more i have a long pretest and you really helped

OpenStudy (johnweldon1993):

Sure hun :)

OpenStudy (anonymous):

thanks wanna message of haha ill post them?

OpenStudy (johnweldon1993):

You can message me if you would like :)